# Environmental impacts of astronomical research infrastructures

### v1

@@ Abstract
Human activities degrade the Earth environment at an unprecedented scale and pace, threatening Earth-system stability, resilience and life-support functions. We can of course deny the facts, get angry about them, try to bargain, or fall into deep depression. Or we may overcome these stages of grief and move towards accepting that human activities need to change, including our own ones. The purpose of this paper is to support astronomers in this transition, by providing insights into the origins of environmental impacts in astronomical research and proposing changes that would make the field sustainable. The paper focuses on the environmental impacts of research infrastructures, since these are the dominant sources of greenhouse gas emissions in astronomy, acknowledging that impact reductions in other areas, for example professional air travelling, need also to be achieved.
@@ 4 Towards sustainable astronomy
Our research suggests that eco-design of astronomical facilities may reduce their environmental impacts, yet these reductions will probably not exceed a few at best. Given the current annual growth rate of in the number of facilities, environmental impact reductions from eco-design will be quickly outpaced by the growth of the field. As we demonstrated, without deliberately limiting the number of facilities and at the same time investing in the decarbonisation of the field, astronomy will not become sustainable. This does not imply that astronomical research will have to stop, nor that no new facilities could be built, but it implies that a more frugal approach needs to be developed.

### v2

@@ Abstract
Human activities degrade the Earth environment at an unprecedented scale and pace, threatening Earth-system stability, resilience and life-support functions. We can of course deny the facts, get angry about them, or try to bargain. Or we may overcome these stages of grief and move towards accepting that human activities need to change, including our own ones. The purpose of this paper is to support astronomers in this transition, by providing insights into the origins of environmental impacts in astronomical research and proposing changes that would make the field sustainable. The paper focuses on the environmental impacts of research infrastructures, since these are the dominant sources of greenhouse gas emissions in astronomy, acknowledging that impact reductions in other areas, for example professional air travelling, need also to be achieved.
@@ 4 Towards sustainable astronomy
Our research suggests that eco-design of astronomical facilities may reduce their environmental impacts, yet these reductions will probably not exceed a few at best. Given the current annual growth rate of 1–3% in the number of facilities, environmental impact reductions from eco-design will be quickly outpaced by the growth of the field. As we demonstrated, without deliberately limiting the number of facilities and at the same time investing in the decarbonisation of the field, astronomy will not become sustainable. This does not imply that astronomical research will have to stop, nor that no new facilities could be built, but it implies that a more frugal approach needs to be developed.